# algorithmic

### v1

@@ Abstract
This work explores multi-modal inference in a high-dimensional simplified model, analytically quantifying the performance gain of multi-modal inference over that of analyzing modalities in isolation. We present the Bayes-optimal performance and weak recovery thresholds in a model where the objective is to recover the latent structures from two noisy data matrices with correlated spikes. The paper derives the approximate message passing (AMP) algorithm for this model and characterizes its performance in the high-dimensional limit via the associated state evolution. The analysis holds for a broad range of priors and noise channels, which can differ across modalities. The linearization of AMP is compared numerically to the widely used partial least squares (PLS) and canonical correlation analysis (CCA) methods, which are both observed to suffer from a sub-optimal recovery threshold.
@@ 1 Introduction
Multi-modal learning in current ML focuses on learning different complex non-linear models of the modalities which ideally cross-inform each other (Ngiam et al., 2011; Baltrušaitis et al., 2018; Bayoudh et al., 2022).
@@
The data model we study is also underlying methods known under the name projection to latent structures (PLS) (Wold, 1975; 1983; Wegelin, 2000), originally referred to as partial least squares (PLS) in the literature, and the more broadly known canonical correlation analysis (CCA) (Hotelling, 1936) subsumed by PLS. These are linear spectral algorithms widely used in chemometrics (Wold et al., 2001; Mehmood et al., 2012), econometrics (Hulland, 1999), neuroscience (Krishnan et al., 2011) and other fields to practically solve linear multi-view inference or prediction tasks in high dimensions.
We provide a typical-case analysis of the Bayes-optimal performance in the high-dimensional limit of the model, based on approximate message passing (AMP) (Donoho et al., 2009; Zdeborová & Krzakala, 2016) with its associated low-dimensional state evolution (SE) (Bayati & Montanari, 2011; Zdeborová & Krzakala, 2016), and the associated Bethe free-energy. This analysis results in the weak recovery threshold that appears as phase transitions in the performance of AMP in the high-dimensional limit. This threshold coincides with the weak-recovery threshold in Bayes-optimal performance if the phase transition is continuous and instead is conjectured to give the optimal performance for polynomial time algorithms in the presence of a first-order transition (Zdeborová & Krzakala, 2016). In the latter case, the Bayes-optimal threshold is determined from the Bethe free energy of the model. We also numerically demonstrate the generally good performance but sub-optimal recovery threshold of PLS even for Gaussian noise channels and priors. For completeness, we show numerical results also for CCA, which is known to have a number of disadvantages compared to "mode-A" PLS (Wegelin, 2000), and is also found here to have less favorable performance and recovery threshold.
@@ 1.1 Spiked Multi-modal Model
where $w^{X}\in\mathbb{R}^{n_{X}}$, $w^{Y}\in\mathbb{R}^{n_{Y}}$ , $v^{X/Y}\in\mathbb{R}^{d}$, and $\xi_{ij}^{X/Y}\overset{\mathrm{iid.}}{\sim}\mathcal{N}(0,\sigma^{2}_{\xi^{X/Y}})$. We assume that $w^{X}_{i}$ and $w^{Y}_{i}$ are independent, while $v^{X}_{j},v^{Y}_{j}$ are given by a correlated joint distribution, such that $X,Y\in\mathbb{R}^{n_{X/Y}\times d}$ are noisy rank-1 matrices with correlated factors $v^{X/Y}$. In the following, the view or modality index is denoted as $z\in\{X,Y\}$ and where needed, the index of the alternate view is denoted as $\bar{z}$. We consider the high-dimensional limit $\lim_{d,n_{z}}\to\infty$ with scaling $\frac{d}{n_{z}}=\alpha_{z}\sim\mathcal{O}(1)$.
@@
$\left.+\frac{1}{2}(R^{z}_{ij}-(S^{z}_{ij})^{2})\frac{\lambda_{z}^{2}}{n_{z}}(w^{z}_{i}v^{z}_{j})^{2}+\mathcal{O}(n^{-\frac{3}{2}}_{z})\right)$
@@
We chose a rank-1 model since we believe it already captures the fundamental phenomenology of the problem.
Note that the signal scales weakly as $n_{z}^{-1/2}$ compared to the $\mathcal{O}(1)$ noise. This is the right scaling to see the BBP transition of the largest singular values correlated to the rank-1 signals disappearing in the random bulk spectra of $X$ and $Y$ at (for unit variances) $\lambda_{z}=\alpha_{z}^{-\frac{1}{4}}$ (Benaych-Georges & Nadakuditi, 2012). We will quantify the improvement that comes from exploiting the correlation between $v^{X}$ and $v^{Y}$ over the BBP thresholds of the two modalities in isolation.
@@ 1.2 Related work
A large number of practical methods for linear multi-view data analysis have been proposed which we do not review in detail. We compare against PLS (Wold, 1975) which exists in several variants (Wegelin, 2000; Rosipal & Krämer, 2006). Notably CCA is equivalent to "mode-B" PLS, but despite its broad popularity is well known for severe shortcomings compared to the canonical "mode-A" PLS (Wegelin, 2000) which we therefore consider instead. These methods are based on the singular value spectrum of the correlation matrix $XY^{T}$ in the case of PLS ("mode-A") and that of the normalized correlation matrix $(XX^{T})^{-\frac{1}{2}}XY^{T}(YY^{T})^{-\frac{1}{2}}$ in the case of CCA.
@@
The canonical PLS algorithm finds rank-k approximations of $X$ and $Y$ by iterating k times the steps: 1) computing the top pair $s_{X},s_{Y}$ of singular vectors of $XY^{T}$, 2) estimating $\hat{v}_{z}=Z^{T}s_{z}$, 3) finding refined estimates $\hat{w}_{z}$ by regressing $Z$ on $\hat{v_{z}}$ so that $\hat{w}_{z}=(\hat{v}_{z}^{T}\hat{v}_{z})^{-1}Z\hat{v}_{z}$, 4) subtracting the rank-1 approximations obtained from each data matrix in isolation, $Z\leftarrow Z-\hat{w}_{z}\hat{v}_{z}^{T}$, 5) repeat from 1). As a simplified variant, PLS-SVD eschews steps 3) and 4), only computing the singular vectors of $XY^{T}$ as the estimates $\hat{w}^{\prime}_{z}=s_{z}$ and again $\hat{v}^{\prime}_{z}=Z^{T}s_{z}$. After the first iteration, which is the only one required in our rank-1 setting, the two variants only differ in that $\hat{w}^{\prime}_{z}=s_{z}$ for PLS-SVD while $\hat{w}_{z}=(\hat{v}_{z}^{T}\hat{v}_{z})^{-1}ZZ^{T}s_{z}$ for PLS-Canonical. The weak recovery thresholds of both variants are thus the same since these estimates only have nonzero overlap with the ground-truth signals $w_{z}$ if the spectrum of $XY^{T}$ has an outlier singular value correlated with the signal.
@@
While the spectrum of $XY^{T}$ has, to our knowledge, not been studied analytically, recent mathematical works exist for the spectrum and BBP-type transition of the normalized correlation matrix in CCA (Bao et al., 2019; Yang, 2022; Bykhovskaya & Gorin, 2023). We show in Figure 3 that the threshold and performance of CCA can be quite far from those of PLS and from the Bayes-optimal values. Empirically, the benefit of shared dimensionality reduction through PLS or CCA compared to single-view methods was analyzed by Abdelaleem et al. (2023), although in a different scaling regime with a stronger signal compared to ours. Non-linear and deep generalizations of CCA have also been developed in the context of self-supervised learning (Balestriero et al., 2023).
@@
The framework we employ is based on a recently matured literature on the statistical physics of algorithmic hardness and Bayes optimal inference (Mézard & Montanari, 2009; Zdeborová & Krzakala, 2016), many aspects of which have now been made rigorous (Bayati & Montanari, 2011; Bolthausen, 2014; Celentano et al., 2020; Krzakala et al., 2023). In particular, we follow largely the notation of Lesieur et al. (2017), who analysed in detail and along related lines a single-view version of the model considered here.
@@
While the single-view spiked matrix model has been studied intensely, e.g. (Rangan & Fletcher, 2012; Lesieur et al., 2017; Montanari & Venkataramanan, 2021), works analysing recovery thresholds for systems that can be related to multi-view or multi-modal learning have so far mainly focused on regression with side information and on variants of community detection. First we note that in mixed matrix-tensor models with rank-1 spike (Sarao Mannelli et al., 2020) the matrix information can be seen as a second view of the rank-1 signal which aids its detection in the tensor data. Kadmon & Ganguli (2018) have applied the AMP framework to low-rank tensor decomposition, where the higher-order tensor can be thought of as data matrices from an experiment with multiple varying conditions forming the additional axes. Compared to our model, this corresponds to more than two views, the rank-1 signals of which only differ by a scalar factor for each additional axis, and no difference in priors is allowed. Rigorous results on AMP for linear regression with side information have been presented by Liu et al. (2019) where the side information is a noisy version of the signal, and by Nandy & Sen (2023) where the side information is given by correlations of signal entries. Chen et al. (2018; 2022) analysed a data matching setting where both views have the same number of features and differ only by their noise realization and a permutation of the feature indices. Deshpande et al. (2018) presented the contextual stochastic block model. Recently an extension was analyzed by Duranthon & Zdeborová (2023), and we note a line of ongoing rigorous work on AMP in multi-view variants of community detection in stochastic block models (Ma & Nandy, 2023; Yang et al., 2024). In these works one view is always a square matrix given by the adjacency matrix of a graph. This is in contrast to our model which can be interpreted as observing an arbitrary number of samples from two views of a correlated latent structure, so that both data matrices are rectangular with $n_{X/Y}$ features and $d$ samples, which allows us to compare with PLS and CCA.
@@ 1.3 Main contributions
  • The information-theoretic performance limits for the multi-view inference task (4), obtained from the state evolution of AMP and the Bethe free-energy.
  • A quantification of the signal-to-noise gain from optimally combining two views, given the prior assumption of a covariance $c_{v}$ between the latent vectors. E.g. for $c_{v}=0.8$ and otherwise unit parameters, recovery is possible from $\sigma_{\xi^{z}}\approx 1.13$, compared to $\sigma_{\xi^{z}}=1$ for the single-view case. We also demonstrate the distance of the recovery threshold of CCA known from Bykhovskaya & Gorin (2023) to the Bayes-optimal value.
  • A spectral method with optimal sensitivity as a linearization of AMP, which combines information from the individual and correlated view, and its comparison to PLS and CCA that both result in sub-optimal sensitivity.
2 Approximate message passing and state evolution
@@ 2.1 AMP
In this section, we discuss the conceptual steps leading from belief propagation (BP) to the AMP algorithm, similarly as undertaken in Lesieur et al. (2017). The technical derivation is given in Appendices A and B. The factor graph of the model is given in Figure 1, corresponding to the BP equations
@@
$n^{z}_{j\to ij}(v^{X}_{j},v^{Y}_{j})$ $=\frac{P_{v}(v^{X}_{j},v^{Y}_{j})}{\mathcal{Z}^{z,n}_{j\to ij}}\prod_{k\neq i}^{n_{z}}\tilde{n}^{z}_{kj\to j}(v^{z}_{j})$
×$\qquad\qquad\quad\times\prod_{k}^{n_{\bar{z}}}\tilde{n}^{\bar{z}}_{kj\to j}(v^{\bar{z}}_{j})$ (8)
@@
where $x\in\mathbb{R}$ for $m^{z}_{i\to ij}(w^{z}_{i})$ and $x\in\mathbb{R}^{2}$ for $n^{z}_{j\to ij}(v^{X}_{J},v^{Y}_{j})$. We also define the normalization of this distribution as $\mathcal{Z}(K,J)=\int\text{d}x\;\mathcal{W}(x,K,J)$, which appears again in the free energy, Section D.3. Interpreting $J$ as a linear source term of the cumulant-generating function $\log\mathcal{Z}(K,J)$ of $x\sim\mathcal{W}(K,J)$, we can write the mean and variance as derivatives w.r.t. the source terms. In compliance with standard notation, we introduce the first derivative (the mean) as the "denoising" function
@@
In the case of $v^{z}$ the off-diagonal terms of $K$ never appear, thus we simplify the notation to $f^{v^{z}}_{\mathrm{in}}(K_{X},K_{Y},J_{X},J_{Y})$ where the $z$ index results from taking the derivative by $J_{X}$ or $J_{Y}$, respectively. However, the term "denoising function" should not obscure the fact that $f^{x}_{\mathrm{in}}(K,J)$ and $\frac{\partial f^{x}_{\mathrm{in}}}{\partial J}(K,J)$ are by definition nothing but the first and second cumulants of the marginal density at the next time step, given by the tilted prior $\mathcal{W}(x,K,J)$.
From rBP (A.14)-(A.17) which is based on the $\mathcal{O}(d^{2})$ messages on the edges of the factor graph, we then obtain AMP which iterates $\mathcal{O}(d)$ node-specific estimates by exploiting that the dependence of the rBP estimates on the target index is weak and can be discounted for by the Onsager reaction term with appropriately delayed time index (Bolthausen, 2014). Concerning the update order determining the time indices of the AMP iteration, while conventionally all messages are passed and updated synchronously for simplicity, there is a freedom to choose an arbitrary update order. Here we choose to update the messages in two sequential blocks, first the marginals of $v^{z}$, then those of $w^{z}$. This is to avoid limit cycles of length 2 arising from the $\mathbb{Z}_{2}$ symmetry in the problem if the relative sign of the $w$ and $v$ estimates does not match. For example, for vanishing noise the otherwise perfect estimate $-w^{z},v^{z}$ would be updated to $w^{z},-v^{z}$ and back to $-w^{z},v^{z}$, etc. The source terms determining the AMP iteration are then
@@
We would like to point out that the time indices $t-1$ for both Onsager reaction terms in (14) and (12) are correct because for the sequential update order, $\hat{v}^{z,t}_{j}$ is updated based on $\hat{w}^{z,t-1}_{i}$ while $\hat{w}^{z,t}_{i}$ is updated based on $\hat{v}^{z,t}_{i}$.
@@
  if "approx. Nishimori" then
@@
  else if "informed" then
@@
  else if "spectral" then

### v2

# Optimal thresholds and algorithms for a model of multi-modal learning in high dimensions

@@ Abstract
This work explores multi-modal inference in a high-dimensional simplified model, analytically quantifying the performance gain of multi-modal inference over that of analyzing modalities in isolation. We present the Bayes-optimal performance and recovery thresholds in a model where the objective is to recover the latent structures from two noisy data matrices with correlated spikes. The paper derives the approximate message passing (AMP) algorithm for this model and characterizes its performance in the high-dimensional limit via the associated state evolution. The analysis holds for a broad range of priors and noise channels, which can differ across modalities. The linearization of AMP is compared numerically to the widely used partial least squares (PLS) and canonical correlation analysis (CCA) methods, which are both observed to suffer from a sub-optimal recovery threshold.
Keywords: Statistical inference, Analysis of algorithms, Machine Learning, Learning theory, Message-passing algorithms, Classical phase transitions
@@ 1 Introduction
Multi-modal learning in current ML focuses on learning different complex non-linear models of the modalities which ideally cross-inform each other [1, 2, 3].
@@
The data model we study is also underlying methods known under the name projection to latent structures (PLS) [4, 5, 6], originally referred to as partial least squares (PLS) in the literature, and the more broadly known canonical correlation analysis (CCA) [7] subsumed by PLS. These are linear spectral algorithms widely used in chemometrics [8, 9], econometrics [10], neuroscience [11] and other fields to practically solve linear multi-view inference or prediction tasks in high dimensions.
We provide a typical-case analysis of the Bayes-optimal performance in the high-dimensional limit of the model, based on approximate message passing (AMP) [12, 13, 14] with its associated low-dimensional state evolution (SE) [15, 13], and the associated Bethe free-energy. This analysis results in the recovery threshold appearing as a phase transition in the performance of AMP in the high-dimensional limit. The threshold of AMP coincides with the Bayes-optimal information theoretic recovery threshold if the phase transition is continuous, and instead is conjectured to give the optimal performance for polynomial time algorithms in the presence of a first-order transition [13]. In the latter case, the Bayes-optimal threshold is determined from the Bethe free energy of the model. We also numerically demonstrate the generally good performance but sub-optimal recovery threshold of PLS even for Gaussian noise channels and priors. We show numerical results also for CCA, which is known to have a number of disadvantages compared to “mode-A” PLS [6], and is also found here to have much less favorable performance and recovery threshold.
@@ 1.1 Spiked Multi-modal Model
where $w^{X}\in\mathbb{R}^{n_{X}}$, $w^{Y}\in\mathbb{R}^{n_{Y}}$ , $v^{X/Y}\in\mathbb{R}^{d}$, and $\xi_{ij}^{X/Y}\overset{\mathrm{iid.}}{\sim}\mathcal{N}(0,\sigma^{2}_{\xi^{X/Y}})$. We assume that $w^{X}_{i}$ and $w^{Y}_{i}$ are independent, while $v^{X}_{j},v^{Y}_{j}$ are given by a correlated joint distribution, such that $X,Y\in\mathbb{R}^{n_{X/Y}\times d}$ are noisy rank-1 matrices with correlated factors $v^{X/Y}$. In the following, the view or modality index is denoted as $z\in\{X,Y\}$ and where needed, the index of the alternate view is denoted as $\bar{z}$. Also we use $Z$ in order to refer to either of the data matrices $X$ or $Y$. We consider the high-dimensional limit $\lim_{d,n_{z}}\to\infty$ with scaling $\frac{d}{n_{z}}=\alpha_{z}\sim\mathcal{O}(1)$. In other words, the aspect ratio of the data matrices $X,Y$ is kept fixed at $\alpha_{X},\alpha_{Y}$ in this so-called proportional limit.
@@
$\qquad\left.+\frac{1}{2}(R^{z}_{ij}-(S^{z}_{ij})^{2})\frac{\lambda_{z}^{2}}{n_{z}}(w^{z}_{i}v^{z}_{j})^{2}+\mathcal{O}(n^{-\frac{3}{2}}_{z})\right)$
@@
We chose a rank-1 model since we believe it already captures the fundamental phenomenology of the problem. An extension to finite rank $r$ would, in analogy to single-view matrix factorization [16, 17], yield an additional index in the equations while the location of the phase transition for the strongest signal direction will not change. Qualitatively different behavior could appear in other scaling limits, e.g. if the signal rank is not finite but proportional to $n_{z}$ and $d$.
Note that the signal scales weakly as $n_{z}^{-1/2}$ compared to the $\mathcal{O}(1)$ noise. This is the right scaling to see the Baik-BenArous-Peché (BBP) transition of the largest singular values correlated to the rank-1 signals disappearing in the random bulk spectra of $X$ and $Y$ at (for unit variances) $\lambda_{z}=\alpha_{z}^{-\frac{1}{4}}$ [18]. We will quantify the improvement that comes from exploiting the correlation between $v^{X}$ and $v^{Y}$ over the BBP thresholds of the two modalities in isolation.
@@ 1.2 Related work
A large number of practical methods for linear multi-view data analysis have been proposed which we do not review in detail. We compare against PLS [4] which exists in several variants [6, 19]. Notably CCA is equivalent to “mode-B” PLS, but despite its broad popularity is well known for severe shortcomings compared to the canonical “mode-A” PLS [6] which we therefore consider instead. These methods are based on the singular value spectrum of the correlation matrix $XY^{T}$ in the case of PLS (“mode-A”) and that of the normalized correlation matrix $(XX^{T})^{-\frac{1}{2}}XY^{T}(YY^{T})^{-\frac{1}{2}}$ in the case of CCA.
@@
The canonical PLS algorithm finds rank-k approximations of $X$ and $Y$ by iterating $k$ times the steps: 1) computing the top pair $s_{X},s_{Y}$ of singular vectors of $XY^{T}$, 2) estimating $\hat{v}_{z}=Z^{T}s_{z}$, 3) finding refined estimates $\hat{w}_{z}$ by regressing $Z$ on $\hat{v}_{z}$ so that $\hat{w}_{z}=(\hat{v}_{z}^{T}\hat{v}_{z})^{-1}Z\hat{v}_{z}$, 4) subtracting the rank-1 approximations obtained from each data matrix in isolation, $Z\leftarrow Z-\hat{w}_{z}\hat{v}_{z}^{T}$, 5) repeat from 1). As a simplified variant, PLS-SVD eschews steps 3) and 4), only computing the singular vectors of $XY^{T}$ as the estimates $\hat{w}^{\prime}_{z}=s_{z}$ and again $\hat{v}^{\prime}_{z}=Z^{T}s_{z}$. After the first iteration, which is the only one required in our rank-1 setting, the two variants only differ in that $\hat{w}^{\prime}_{z}=s_{z}$ for PLS-SVD while $\hat{w}_{z}=(\hat{v}_{z}^{T}\hat{v}_{z})^{-1}ZZ^{T}s_{z}$ for PLS-Canonical. The recovery thresholds of both variants are thus the same since these estimates only have nonzero overlap with the ground-truth signals $w_{z}$ if the spectrum of $XY^{T}$ has an outlier singular value correlated with the signal.
@@
While the spectrum of $XY^{T}$ has, to our knowledge, not been studied analytically, recent mathematical works exist for the spectrum and BBP-type transition of the normalized correlation matrix in CCA [20, 21, 22]. We show in Figure 4 that the threshold and performance of CCA can be quite far from those of PLS and from the Bayes-optimal values. Empirically, the benefit of shared dimensionality reduction through PLS or CCA compared to single-view methods was analyzed in [23], although in a different scaling regime with a stronger signal compared to ours. Non-linear and deep generalizations of CCA have also been developed in the context of self-supervised learning [24].
@@
The framework we employ is based on a recently matured literature on the statistical physics of algorithmic hardness and Bayes optimal inference [25, 13], many aspects of which have now been made rigorous [15, 26, 27, 28, 14]. In particular, we follow largely the notation of [17], who analyzed in detail and along related lines a single-view version of the model considered here. For a physics focused introduction to AMP and algorithmic hardness we refer the reader to [13], for a mathematical introduction to [14].
@@
While the single-view spiked matrix model has been studied intensely, e.g. [16, 17, 29], works analysing recovery thresholds for systems that can be related to multi-view or multi-modal learning have so far mainly focused on regression with side information and on variants of community detection. First we note that in mixed matrix-tensor models with rank-1 spike [30] the matrix information can be seen as a second view of the rank-1 signal which aids its detection in the tensor data. [31] applied the AMP framework to low-rank tensor decomposition, where the higher-order tensor can be thought of as data matrices from an experiment with multiple varying conditions forming the additional axes. Compared to our model, this corresponds to more than two views, the rank-1 signals of which only differ by a scalar factor for each additional axis, and no difference in priors is allowed. Rigorous results on AMP for linear regression with side information have been presented in [32] where the side information is a noisy version of the signal, and in [33] where the side information is given by correlations of signal entries. [34, 35] analyzed a data matching setting where both views have the same number of features and differ only by their noise realization and a permutation of the feature indices. [36] presented the contextual stochastic block model, of which recently an extension was analyzed in [37].
Finally, we note two lines of closely related mathematical works. AMP in multi-view variants of community detection in stochastic block models was developed in parallel to our study in [38, 39] and has shortly after the initial publication of our manuscript resulted in “OrchAMP” [40] where a multi-view inference model similar to ours is considered which also allows for low-dimensional side information. In the latter work Nandy and Ma focus on developing a practical algorithm for sufficiently strong signals (above the BBP threshold for each single view) but unknown priors, using iterative updating of the denoising function through the empirical Bayes prior [41], and validate their method on cell atlas construction; while we here focus on the improvement of Bayes-optimal recovery limits over the single view BBP threshold, and their comparison to the performance of CCA and PLS. During the review process, additionally work on the information-theoretic inference limits [42] and recently AMP [43] of a very general matrix-tensor product model has been brought to our attention. This meta-model subsumes a variety of matrix models with block-structured correlations, the stochastic block model, and tensor-data settings where only the first two index dimensions are extensive. Also our reductionist multi-view model (1),(2) can be brought into such a matrix-tensor product form. Although the authors of [42, 43] did not consider a corresponding setting of priors it would be possible to derive our AMP and state evolution starting from their framework. The analysis of the matrix-tensor product model illustrates that it is possible to derive AMP and abstract state evolution iterations at a more general level, therefore our derivation may rightly be seen as a concrete exposition of these by now established methods in the language of statistical physics, rather than a methodological advance.
@@ 1.3 Main contributions
  • The information-theoretic and algorithmic performance limits for the multi-view inference task (4), obtained from the state evolution of AMP and the Bethe free-energy (Section 3). These results could also be obtained as a consequence of the more generic MTP framework from [43].
  • A quantification of the signal-to-noise gain from optimally combining two views, given the prior assumption of a covariance $c_{v}$ between the latent vectors (Section 3.2 and Figure 2). E.g. for $c_{v}=0.8$ and otherwise unit parameters, recovery is possible from $\sigma_{\xi^{z}}\approx 1.13$, compared to $\sigma_{\xi^{z}}=1$ for the single-view case. We also demonstrate the distance of the recovery threshold of CCA known from [22] to the Bayes-optimal value (Figure 4).
  • A spectral method as a linearization of AMP, which therefore has the same optimal recovery threshold as AMP and combines information from the individual and correlated view (Section 2.2), and a numerical comparison to PLS and CCA that both result in sub-optimal sensitivity (Section 4). In particular, PLS becomes sub-optimal when the correlation between the latent vectors is small (Figure 5), and CCA generally suffers from lower sample efficiency (Figure 4).
A reader interested mainly in the results may safely skip the methods Section 2 and continue from the terminology primer Section 3.1 or the main results starting from Section 3.2.
2 Methods: Approximate message passing and state evolution
@@ 2.1 AMP
In this section, we discuss the conceptual steps leading from belief propagation (BP) to the AMP algorithm. The main technical contributions here are the formulation of a parsimonious multi-view model in Section 1.1 and the treatment of correlated latent variables by two-dimensional marginals in the BP messages. The remaining derivation of AMP and the state evolution then goes through as a straight-forward generalization of the calculations for single-view matrix factorization presented by [17], whose notation we adapt slightly for more consistency with standard symbols in statistical physics. The technical derivation is given in A and B. The factor graph of the model is given in Figure 1, corresponding to the BP equations
@@
$n^{z}_{j\to ij}(v^{X}_{j},v^{Y}_{j})$ $=\frac{P_{v}(v^{X}_{j},v^{Y}_{j})}{\mathcal{Z}^{z,n}_{j\to ij}}\prod_{k\neq i}^{n_{z}}\tilde{n}^{z}_{kj\to j}(v^{z}_{j})\prod_{k}^{n_{\bar{z}}}\tilde{n}^{\bar{z}}_{kj\to j}(v^{\bar{z}}_{j})$ (8)
@@
where $x\in\mathbb{R}$ for $m^{z}_{i\to ij}(w^{z}_{i})$ and $x\in\mathbb{R}^{2}$ for $n^{z}_{j\to ij}(v^{X}_{J},v^{Y}_{j})$. We also define the normalization of this distribution as $\mathcal{Z}(K,J)=\int\text{d}x\;\mathcal{W}(x,K,J)$, which appears again in the free energy, Section D.3. Interpreting $J$ as a linear source term of the cumulant-generating function $\log\mathcal{Z}(K,J)$ of $x\sim\mathcal{W}(K,J)$, we can write the mean and variance as derivatives w.r.t. the source terms. In compliance with standard notation, we introduce the first derivative (the mean) as the “denoising” function
@@
In the case of $v^{z}$ the off-diagonal terms of $K$ never appear, thus we simplify the notation to $f^{v^{z}}_{\mathrm{in}}(K_{X},K_{Y},J_{X},J_{Y})$ where the $z$ index results from taking the derivative by $J_{X}$ or $J_{Y}$, respectively. However, the term “denoising function” should not obscure the fact that $f^{x}_{\mathrm{in}}(K,J)$ and $\frac{\partial f^{x}_{\mathrm{in}}}{\partial J}(K,J)$ are by definition nothing but the first and second cumulants of the marginal density at the next time step, given by the tilted prior $\mathcal{W}(x,K,J)$.
From rBP (45)-(48) which is based on the $\mathcal{O}(d^{2})$ messages on the edges of the factor graph, we then obtain AMP which iterates $\mathcal{O}(d)$ node-specific estimates by exploiting that the dependence of the rBP estimates on the target index is weak and can be discounted for by a so-called cavity- or Onsager-reaction term with appropriately delayed time index [26]. That this correction holds in the asymptotic limit for factor graphs of the structure considered here is a consequence of Theorem 1 of [43]. Concerning the update order determining the time indices of the AMP iteration, while conventionally all messages are passed and updated synchronously for simplicity, there is a freedom to choose an arbitrary update order. Here we choose to update the messages in two sequential blocks, first the marginals of $v^{z}$, then those of $w^{z}$. This is to avoid limit cycles of length 2 arising from the $\mathbb{Z}_{2}$ symmetry in the problem if the relative sign of the $w$ and $v$ estimates does not match. For example, for vanishing noise the otherwise perfect estimate $w^{z},v^{z}$ would be updated to $w^{z},-v^{z}$ and back to $w^{z},v^{z}$, etc. The source terms determining the AMP iteration (Algorithm 1) are then
@@
We would like to point out that the time indices t1t-1 for both Onsager reaction terms in (14) and (12) are correct because for the sequential update order, $\hat{v}^{z,t}_{j}$ is updated based on $\hat{w}^{z,t-1}_{i}$ while $\hat{w}^{z,t}_{i}$ is updated based on $\hat{v}^{z,t}_{i}$. Again the form of the AMP iteration can also be specialized from the generic form of [43]. We will now explore the concrete properties of the algorithm for the model considered here.
@@
  init_method ’spectral’/’informed’/’approx. Nishimori’
@@
  # See B.3 for a discussion of initializations.
if init_method == ’approx. Nishimori’ then
@@
else if init_method == ’informed’ then
@@
else if init_method == ’spectral’ then